OpenStudy (anonymous):

pleasee help! Solve. x2 - 8x = 20 What are the solutions of 4x2 - 3x - 5 = 0?

OpenStudy (anonymous):

Simplifying x2 + -8x = 20 Reorder the terms: -8x + x2 = 20 Solving -8x + x2 = 20 Solving for variable 'x'. Reorder the terms: -20 + -8x + x2 = 20 + -20 Combine like terms: 20 + -20 = 0 -20 + -8x + x2 = 0 Factor a trinomial. (-2 + -1x)(10 + -1x) = 0 Subproblem 1Set the factor '(-2 + -1x)' equal to zero and attempt to solve: Simplifying -2 + -1x = 0 Solving -2 + -1x = 0 Move all terms containing x to the left, all other terms to the right. Add '2' to each side of the equation. -2 + 2 + -1x = 0 + 2 Combine like terms: -2 + 2 = 0 0 + -1x = 0 + 2 -1x = 0 + 2 Combine like terms: 0 + 2 = 2 -1x = 2 Divide each side by '-1'. x = -2 Simplifying x = -2 Subproblem 2Set the factor '(10 + -1x)' equal to zero and attempt to solve: Simplifying 10 + -1x = 0 Solving 10 + -1x = 0 Move all terms containing x to the left, all other terms to the right. Add '-10' to each side of the equation. 10 + -10 + -1x = 0 + -10 Combine like terms: 10 + -10 = 0 0 + -1x = 0 + -10 -1x = 0 + -10 Combine like terms: 0 + -10 = -10 -1x = -10 Divide each side by '-1'. x = 10 Simplifying x = 10Solutionx = {-2, 10}

OpenStudy (anonymous):

Simplifying 4x2 + -3x + -5 = 0 Reorder the terms: -5 + -3x + 4x2 = 0 Solving -5 + -3x + 4x2 = 0 Solving for variable 'x'. Begin completing the square. Divide all terms by 4 the coefficient of the squared term: Divide each side by '4'. -1.25 + -0.75x + x2 = 0 Move the constant term to the right: Add '1.25' to each side of the equation. -1.25 + -0.75x + 1.25 + x2 = 0 + 1.25 Reorder the terms: -1.25 + 1.25 + -0.75x + x2 = 0 + 1.25 Combine like terms: -1.25 + 1.25 = 0.00 0.00 + -0.75x + x2 = 0 + 1.25 -0.75x + x2 = 0 + 1.25 Combine like terms: 0 + 1.25 = 1.25 -0.75x + x2 = 1.25 The x term is -0.75x. Take half its coefficient (-0.375). Square it (0.140625) and add it to both sides. Add '0.140625' to each side of the equation. -0.75x + 0.140625 + x2 = 1.25 + 0.140625 Reorder the terms: 0.140625 + -0.75x + x2 = 1.25 + 0.140625 Combine like terms: 1.25 + 0.140625 = 1.390625 0.140625 + -0.75x + x2 = 1.390625 Factor a perfect square on the left side: (x + -0.375)(x + -0.375) = 1.390625 Calculate the square root of the right side: 1.179247642 Break this problem into two subproblems by setting (x + -0.375) equal to 1.179247642 and -1.179247642. Subproblem 1x + -0.375 = 1.179247642 Simplifying x + -0.375 = 1.179247642 Reorder the terms: -0.375 + x = 1.179247642 Solving -0.375 + x = 1.179247642 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '0.375' to each side of the equation. -0.375 + 0.375 + x = 1.179247642 + 0.375 Combine like terms: -0.375 + 0.375 = 0.000 0.000 + x = 1.179247642 + 0.375 x = 1.179247642 + 0.375 Combine like terms: 1.179247642 + 0.375 = 1.554247642 x = 1.554247642 Simplifying x = 1.554247642 Subproblem 2x + -0.375 = -1.179247642 Simplifying x + -0.375 = -1.179247642 Reorder the terms: -0.375 + x = -1.179247642 Solving -0.375 + x = -1.179247642 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '0.375' to each side of the equation. -0.375 + 0.375 + x = -1.179247642 + 0.375 Combine like terms: -0.375 + 0.375 = 0.000 0.000 + x = -1.179247642 + 0.375 x = -1.179247642 + 0.375 Combine like terms: -1.179247642 + 0.375 = -0.804247642 x = -0.804247642 Simplifying x = -0.804247642 SolutionThe solution to the problem is based on the solutions from the subproblems. x = {1.554247642, -0.804247642}

OpenStudy (anonymous):

thanks so much you are a lifesaver!